OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)2 – 25? Identify the x–intercepts by graphing the quadratic equation i need the answer

OpenStudy (anonymous):

when u simplify that.... it results in y=2x-31----- assuming that is in y=mx+b form, where mx= the slope ans b= y intercept 2/1 is the slope and -31= y intercept